# Impact of Fixing Spins in a Quantum Annealer with Energy Rescaling

### v1

@@ 4 Conclusion
We investigated the effects of rescaling and fixing spins through both experimental hardware and numerical simulations. The experiment clarifies the impact of fixing spins and rescaling on a homogeneous fully connected ferromagnetic Ising model. Our results demonstrate that fixing spins effectively leads to a low energy state in quantum annealing hardware. Additionally, fixing spins preserves the spin chain generated by the minor embedding, which is an inherent feature when embedding the Ising model into hardware.
These findings suggest that the size-reduction method requires a broader range of parameters to address larger-scale problems on hardware. Furthermore, to solve large-scale problems effectively, a fixing spins method with a low error probability is essential. The results also emphasize that increasing the number of couplers for each qubit is a powerful strategy to prevent chain breaking, as avoiding long spin chains helps mitigate the chain strength issue.
@@
This study investigated the simplest setup to explore the complex relationships between fixing spins and rescaling. Future work will involve investigating other models, such as the random Ising model and combinatorial optimization with constraints, which exhibit different properties of local magnetic fields that increase after fixing spins.
@@ Appendix A Energy-gap analysis by bosonization in thermodynamic limit
$\mathcal{H}(t)$ $=(1-s)\mathcal{H}_{\mathrm{d}}+s\mathcal{H}_{\mathrm{p}}$
$=(1-s)\sum_{i=1}^{N}\sigma_{i}^{x}+s\left(-\frac{h}{r}\sum_{i=1}^{N}\sigma_{i}^{z}-\frac{J}{rN}\sum_{1\leq i<j\leq N}\sigma_{i}^{z}\sigma_{j}^{z}\right)$
$=(1-s)S_{x}-s\frac{h}{r}S_{z}-\frac{sJ}{rN}(S_{z})^{2},$ (16)
where $S_{x}$ and $S_{z}$ are Pauli matrices in the composite system. Considering a rotation around the $y$-axis by an angle $\theta_{0}$, the composite spins can be expressed as follows:
@@
$\mathcal{H}(s)$ $=(1-s)(-\sin{\theta_{0}}\tilde{S}_{x}+\cos{\theta_{0}}\tilde{S}_{z})$
$-s\frac{h}{r}(\cos{\theta_{0}}\tilde{S}_{x}+\sin{\theta_{0}}\tilde{S}_{z})-\frac{sJ}{rN}(\cos{\theta_{0}}\tilde{S}_{x}+\sin{\theta_{0}}\tilde{S}_{z})^{2}.$ (27)
After applying the Holstein–Primacoff transformation[42] to Eq. (27), the total Hamiltonian is expressed in terms of the creation and annihilation operator $a,a^{{\dagger}}$. The Holstein–Primacoff transformation is given by:
$\tilde{S}_{z}$ $=\frac{N}{2}-aa^{{\dagger}},$ (28)
@@
$\tilde{S}_{x}$ $=\frac{1}{2}\sqrt{N-a^{{\dagger}}a}(a+a^{{\dagger}})=\frac{\sqrt{N}}{2}\sqrt{1-\frac{a^{{\dagger}}a}{N}}(a+a^{{\dagger}}).$ (31)
@@
$\tilde{S}_{z}$ $=\frac{N}{2}-aa^{{\dagger}},$ (32)
@@
$\mathcal{H}(s)=$ $\left[-s\frac{h}{r}\sin{\theta_{0}}-(1-s)\cos{\theta_{0}}-\frac{sJ\sin^{2}{\theta_{0}}}{2r}\right]\frac{N}{2}$
$-\left[s\frac{h}{r}\cos{\theta_{0}}+(1-s)\sin{\theta_{0}}+\right.$
$+\left.s\frac{J}{r}\cos{\theta_{0}}\sin{\theta_{0}}\right](a+a^{{\dagger}})\frac{\sqrt{N}}{2}$
$+\left[s\frac{h}{r}\sin{\theta_{0}}+(1-s)\cos{\theta_{0}}\right]a^{{\dagger}}a$
$-\frac{sJ}{4r}\cos^{2}{\theta_{0}}[a^{2}+aa^{{\dagger}}+a^{{\dagger}}a+(a^{{\dagger}})^{2}]+s\frac{J}{r}\sin^{2}{\theta_{0}}a^{{\dagger}}a$
$+\frac{sJ}{2r\sqrt{N}}\cos{\theta_{0}}\sin{\theta_{0}}a^{{\dagger}}a(a+a^{{\dagger}})-\frac{sJ}{rN}\sin^{2}{\theta_{0}}(a^{{\dagger}}a)^{2}.$ (34)
@@
$\mathcal{H}(s)\approx$ $\left[-\frac{h}{r}\sin{\theta_{0}}-(1-s)\cos{\theta_{0}}-\frac{sJ\sin^{2}{\theta_{0}}}{2r}\right]\frac{N}{2}$
$-\left[s\frac{h}{r}\cos{\theta_{0}}+(1-s)\sin{\theta_{0}}\right.$
$+\left.s\frac{J}{r}\cos{\theta_{0}}\sin{\theta_{0}}\right](a+a^{{\dagger}})\frac{\sqrt{N}}{2}$
$+\left[s\frac{h}{r}\sin{\theta_{0}}+(1-s)\cos{\theta_{0}}\right.$
$+\left.s\frac{J}{r}\sin^{2}{\theta_{0}}-\frac{sJ}{2r}\cos^{2}{\theta_{0}}\right]a^{{\dagger}}a$
$-\frac{sJ}{4r}\cos^{2}{\theta_{0}}[a^{2}+(a^{{\dagger}})^{2}]-\frac{sJ}{4r}\cos^{2}{\theta_{0}}.$ (35)
@@
$e\equiv$ $\frac{1}{2}\left[-s\frac{h}{r}\sin{\theta_{0}}-(1-s)\cos{\theta_{0}}-\frac{sJ\sin^{2}{\theta_{0}}}{2r}\right],$ (36)
$\delta\equiv$ $\left[s\frac{h}{r}\sin{\theta_{0}}+(1-s)\cos{\theta_{0}}+s\frac{J}{r}\sin^{2}{\theta_{0}}-\frac{sJ}{2r}\cos^{2}{\theta_{0}}\right],$ (37)
$\gamma\equiv$ $\frac{sJ}{4r}\cos^{2}{\theta_{0}}.$ (38)
@@
$\frac{\partial e}{\partial\theta_{0}}=-s\frac{h}{r}\cos{\theta_{0}}+(1-s)\sin{\theta_{0}}-s\frac{J}{r}\cos{\theta_{0}}\sin{\theta_{0}}=0.$ (39)
@@
$\Delta$ $=\delta\sqrt{1-\epsilon^{2}}$
=$=\left\{s^{2}\sin^{2}{\theta_{0}}\left(\frac{h}{r}+\frac{1+\sin^{2}{\theta_{0}}}{s\sin{\theta_{0}}}\right)^{2}-\frac{(1+\sin^{2}{\theta_{0}})^{2}}{s^{2}\sin^{2}{\theta_{0}}}\right.$
$+\left.s^{2}\sin^{4}{\theta_{0}}\frac{J^{2}}{r^{2}}+s[2\sin^{2}{\theta_{0}}-(1-s)\cos^{3}{\theta_{0}}]\frac{J}{r}\right.$
$\left.+(1-s)^{2}\cos^{2}{\theta_{0}}\right\}^{\frac{1}{2}}.$ (45)

### v3

@@ 4 Conclusion
We investigated the effects of rescaling and fixing spins through experimental hardware and numerical simulations. The experiment clarifies the impact of fixing spins and rescaling on a homogeneous fully connected ferromagnetic Ising model. Our results demonstrate that fixing spins effectively leads to a low-energy state in quantum annealing hardware. Additionally, fixing spins preserves the qubit chain generated by the minor embedding, which is an inherent feature when embedding the Ising model into hardware.
These findings suggest that the size-reduction method requires a broader range of parameters to address larger-scale problems on hardware. Furthermore, to solve large-scale problems effectively, a fixing spins method with a low error probability is essential. However, excessive reduction in problem size can adversely affect the energy scale, potentially compromising solution quality or hardware compatibility.
Based on these results, we can make proposals for hardware implementation. The results emphasize that increasing the number of couplers for each qubit is a powerful strategy to prevent chain break. Increasing the number of couplers shortens the chain length. When the chain length is reduced, the local magnetic fields at each qubit become stronger as the distribution of the local magnetic fields across the qubits in a qubit chain is suppressed. For the same range of magnetic fields and interactions in the Ising model, the ratio of the local magnetic fields to chain strength becomes larger, making chain breaks less likely. Therefore, increasing the number of couplers helps prevent chain breaks.
@@
This study investigated the simplest setup to explore the complex relationships between fixing spins and rescaling. Future work will investigate other models, such as the random Ising model and combinatorial optimization with constraints, which exhibit different properties of local magnetic fields that increase after fixing spins.
@@ Appendix A Energy-gap analysis by bosonization in thermodynamic limit
$\mathcal{H}(s)$ $=(1-s)\mathcal{H}_{\mathrm{d}}+s\mathcal{H}_{\mathrm{p}}$
$=-(1-s)\sum_{i=1}^{N}\sigma_{i}^{x}+s\left(-\frac{h}{r}\sum_{i=1}^{N}\sigma_{i}^{z}-\frac{J}{rN}\sum_{1\leq i<j\leq N}\sigma_{i}^{z}\sigma_{j}^{z}\right)$
$=-(1-s)2S_{x}-s\frac{h}{r}2S_{z}-\frac{sJ}{2rN}\left(2S_{z}\right)^{2}+\mathrm{const},$ (16)
where $S_{x}$ and $S_{z}$ are $x$-component and $z$-component of the total spin variable in the composite system, respectively. Considering a rotation around the $y$-axis by an angle $\theta_{0}$, the composite spins can be expressed as follows:
@@
$\mathcal{H}(s)$ $=-2(1-s)(-\sin{\theta_{0}}\tilde{S}_{x}+\cos{\theta_{0}}\tilde{S}_{z})$
$-s\frac{2h}{r}(\cos{\theta_{0}}\tilde{S}_{x}+\sin{\theta_{0}}\tilde{S}_{z})-\frac{2sJ}{rN}(\cos{\theta_{0}}\tilde{S}_{x}+\sin{\theta_{0}}\tilde{S}_{z})^{2}.$ (27)
After applying the Holstein–Primacoff transformation[43] to Eq. (27), the total Hamiltonian is expressed in terms of the creation and annihilation operators $a,a^{{\dagger}}$. The Holstein–Primacoff transformation is given by:
$\tilde{S}_{z}$ $=\frac{N}{2}-a^{{\dagger}}a,$ (28)
@@
$\tilde{S}_{x}$ $=\tilde{S}_{+}+\tilde{S}_{-}$
$=\frac{\sqrt{N}}{2}\left(\sqrt{1-\frac{a^{\dagger}a}{N}}a+a^{\dagger}\sqrt{1-\frac{a^{\dagger}a}{N}}\right).$ (31)
@@
$\tilde{S}_{z}$ $=\frac{N}{2}-a^{{\dagger}}a,$ (32)
@@
$\mathcal{H}(s)=$ $\left[2(1-s)\cos{\theta_{0}}-s\frac{2h}{r}\sin{\theta_{0}}-\frac{sJ\sin^{2}{\theta_{0}}}{r}\right]\frac{N}{2}$
$+\left[2(1-s)\sin{\theta_{0}}-s\frac{2h}{r}\cos{\theta_{0}}\right.$
$-\left.s\frac{2J}{r}\cos{\theta_{0}}\sin{\theta_{0}}\right](a+a^{{\dagger}})\frac{\sqrt{N}}{2}+2(1-s)\cos\theta_{0}$
$+s\frac{2h}{r}\sin{\theta_{0}}-s\frac{J}{r}\cos^{2}\theta_{0}+s\frac{2J}{r}\sin^{2}{\theta_{0}}a^{\dagger}a$
$-s\frac{J}{2r}\cos^{2}\theta_{0}\left[(a^{\dagger})^{2}+a^{2}+1\right]$
$-\frac{sJ}{r\sqrt{N}}\cos{\theta_{0}}\sin{\theta_{0}}[a^{{\dagger}}a(a+a^{{\dagger}})+(a^{\dagger}+a)a^{\dagger}a]$
$-s\frac{2J}{2rN}\sin^{2}{\theta_{0}}(a^{{\dagger}}a)^{2}.$ (34)
@@
$\mathcal{H}(s)\approx$ $\left[2(1-s)\cos{\theta_{0}}-s\frac{2h}{r}\sin{\theta_{0}}-\frac{sJ\sin^{2}{\theta_{0}}}{r}\right]\frac{N}{2}$
$+\left[2(1-s)\sin{\theta_{0}}-s\frac{2h}{r}\cos{\theta_{0}}\right.$
$-\left.s\frac{2J}{r}\cos{\theta_{0}}\sin{\theta_{0}}\right](a+a^{{\dagger}})\frac{\sqrt{N}}{2}+2(1-s)\cos\theta_{0}$
$+s\frac{2h}{r}\sin{\theta_{0}}-s\frac{J}{r}\cos^{2}\theta_{0}+s\frac{2J}{r}\sin^{2}{\theta_{0}}a^{\dagger}a$
$-s\frac{J}{2r}\cos^{2}\theta_{0}\left[(a^{\dagger})^{2}+a^{2}+1\right].$ (35)
@@
$e\equiv$ $\frac{1}{2}\left[-2(1-s)\cos{\theta_{0}}-s\frac{2h}{r}\sin{\theta_{0}}-s\frac{J}{r}\sin^{2}{\theta_{0}}\right],$ (36)
$\delta\equiv$ $2(1-s)\cos{\theta_{0}}+s\frac{2h}{r}\sin{\theta_{0}}+s\frac{2J}{r}\sin^{2}{\theta_{0}}-s\frac{J}{r}\cos^{2}{\theta_{0}},$ (37)
$\gamma\equiv$ $-\frac{sJ}{2r}\cos^{2}{\theta_{0}}.$ (38)
@@
$\frac{\partial e}{\partial\theta_{0}}=\frac{1}{2}\left[2(1-s)\sin{\theta_{0}}-s\frac{2h}{r}\cos{\theta_{0}}-s\frac{2J}{r}\cos{\theta_{0}}\sin{\theta_{0}}\right]=0.$ (39)
@@
$\Delta(s)$ $=\delta\sqrt{1-\epsilon^{2}}$
=$=\left\{s^{2}\sin^{2}{\theta_{0}}\left(\frac{2h}{r}+\frac{1+\sin^{2}{\theta_{0}}}{s\sin{\theta_{0}}}\right)^{2}-\frac{(1+\sin^{2}{\theta_{0}})^{2}}{s^{2}\sin^{2}{\theta_{0}}}\right.$
$+\left.s^{2}\sin^{4}{\theta_{0}}\frac{4J^{2}}{r^{2}}+s[2\sin^{2}{\theta_{0}}-2(1-s)\cos^{3}{\theta_{0}}]\frac{2J}{r}\right.$
$\left.+4(1-s)^{2}\cos^{2}{\theta_{0}}\right\}^{\frac{1}{2}},\quad 0\leq s\leq 1.$ (45)